# SpecOffload: Unlocking Latent GPU Capacity for LLM Inference on Resource-Constrained Devices

### v1

@@ Abstract
Efficient LLM inference on resource-constrained devices presents significant challenges in compute and memory utilization. Due to limited GPU memory, existing systems offload model weights to CPU memory, incurring substantial I/O overhead between the CPU and GPU. This leads to two major inefficiencies: (1) GPU cores are underutilized, often remaining idle while waiting for data to be loaded; and (2) GPU memory has low impact on performance, as reducing its capacity has minimal effect on overall throughput. In this paper, we propose SpecOffload, a high-throughput inference engine that embeds speculative decoding into offloading. Our key idea is to unlock latent GPU resources for storing and executing a draft model used for speculative decoding, thus accelerating inference at near-zero additional cost. To support this, we carefully orchestrate the interleaved execution of target and draft models in speculative decoding within the offloading pipeline, and propose a planner to manage tensor placement and select optimal parameters. Compared to the best baseline, SpecOffload improves GPU core utilization by 4.49× and boosts inference throughput by 2.54×. Our code is available at https://github.com/MobiSense/SpecOffload.
@@ 1 Introduction
As Large Language Models (LLMs) evolve, their real-world use extends far beyond chatbots to diverse applications including synthetic data generation datageneration , form processing formprocessing , and data wrangling datawrangling . These tasks are characterised by LLMs conducting offline inference in batches over large number of tokens. For instance, corporations need to process all archives of financial documentation, whilst individuals want to construct knowledge repositories from accumulated materials. In such workloads, higher inference throughput (the number of tokens generated divided by total generation time, token/s) translates into lower total completion time, hence is the key metric.
Privacy and cost concerns drive these tasks toward LLM deployment on edge servers or PCs, where GPU memory constrains. For instance, conducting inference exclusively using GPU memory for Mixtral 8x22B (282GB) Mixtral8x22b requires at least four NVIDIA H100 (80GB) GPUs. Offloading is one of the mainstream solutions to memory-constrained inference, transferring most model parameters to more economical, capacious CPU memory and reloading them to GPU memory only when computation demands. There are also methods to overcome the memory bottleneck by compressing the model and KV cache, such as quantization, pruning, sparsification quantization ; pruning ; sparsification , etc., which are orthogonal and can be applied on top of offloading. Our focus is designing efficient offloading strategies for high-throughput inference on resource-constrained devices.
@@
  • Underutilization of GPU cores. As shown in Figure 2, during the decoding phase, the average GPU core utilization of existing methods is only 13% at most. This inefficiency stems from GPU cores frequently remaining long time idle while awaiting parameters loading. To alleviate this issue, existing methods increase the batch size to amortize I/O overhead—by loading model parameters once and reusing them across the enlarged batch, thereby improving throughput Flexgen ; Klotski . However, due to the limitation of GPU memory capacity or CPU computational capabilities, scalability of the batch size remains inherently limited. The maximum batch size achieved by the SOTA scheme in Figure 2 is only 64, insufficient to bridge the huge gap between I/O and GPU latency (even with this batch size, the gap remains over 10×).
  • Marginal utility of GPU memory. When models far exceed GPU memory capacity, throughput remains almost unchanged even when memory usage is greatly reduced during decoding phase. We conducted experiments with FlexGen Flexgen , the most effective solution among SOTA approaches. FlexGen offloads attention computations to CPU while computing FFN layers on GPU during the decoding phase. As illustrated in Figure 2, reducing memory usage by over 5.42× for Mixtral 8×7B models led to only a 13% drop in throughput; similarly, a 2.89× reduction for Mixtral 8×22B models caused just a 5% decline. This phenomenon arises because the total volume of data to be loaded remains nearly constant, as the model size vastly exceeds the available GPU memory—leaving little room for variation even when accounting for partical parameters that can reside permanently in memory. For instance, generating a single token necessitates loading all (56) FFN layers for the leftmost blue point in Figure 2, compared to 53 layers for the rightmost blue point. Consequently, memory reduction precipitates only negligible changes in total inference time and throughput.
To harnesses GPU compute and memory resources more efficiently, we design SpecOffload, a novel offloading framework that unlocks latent GPU capacity by leveraging speculative decoding (SD). SD is a technique that accelerates generation by employing an auxiliary lightweight draft model to produce multiple candidate tokens, which are subsequently verified in parallel by the target model, enabling the generation of multiple tokens per forward pass blockwise ; FastInference ; ACloseLook .
@@
  • Computing draft model during GPU core idleness: SD requires draft model to generate multiple candidate tokens in advance. Given the substantial idle periods prevalent in existing offloading frameworks, these intervals can be utilized for completing the draft model’s computational tasks.
  • Storing draft model use "low-yield" GPU memroy: SD requires loading a draft model into memory for draft generation. We can repurpose "low-yield" memory allocations from existing frameworks to store draft model parameters and its caches instead. For instance, as shown in Figure 2, extracting 17GB of "low-yield" memory allocation provides sufficient capacity for a draft model such as Mistral 7B mistral7b to operate normally within the GPU at a small batch.
@@
  • We conduct a quantitative analysis of GPU resource utilization in representative scenarios and identify key limitations in SOTA frameworks——underutilization of GPU cores and marginal utility of GPU memory, thus reveal a novel perspective for enhancing offloading performance.
  • By designing sub-layer model decomposition and fine-grained scheduling compute and memory resources, we delicately embedded SD into offloading with virtually zero overhead, thereby increasing GPU core utilization by 4.49 times.
@@ 2.1 Generative LLM Inference
Large language models consist of stacked Transformer layers attentionisallyouneed . During inference, tokens are generated in an autoregressive paradigm: the prefill phase processes the complete input sequence to construct the KV cache, whereas in the decoding phase, subsequent token is produced based on the previously generated token and the cached KV states pageattention ; pope_efficiently_2023 ; GPT3 .
@@ 2.2 Speculative Decoding
Speculative decoding (SD) is a method for accelerating LLM inference. It adheres to a “Draft-then-Verify” framework: at each decoding step, a lightweight draft model initially proposes multiple candidate tokens (e.g., $(\hat{w}_{1},\hat{w}_{2},\hat{w}_{3},\hat{w}_{4}$)), which are collectively verified by a larger target model in a single forward pass. Only the valid subset $({w}_{1},{w}_{2})$ is accepted, after which the target model resumes decoding by independently generating the subsequent token $w_{3}$ blockwise ; FastInference ; ACloseLook . This approach enables the target model to generate multiple tokens per inference step. To further enhance the efficiency of SD, prior research has predominantly explored two avenues: the design of more effective draft models distillspec ; draft-verify ; onlineSD and draft structures medusa ; specinfer ; specexec . However, traditional single-batch SD is not well-suited for integration with offloading, as the computations of the draft and target models must be executed sequentially. As a result, the GPU resources still remain underutilized during the target model’s verification phase. By introducing a dual-batch rotation strategy, SpecOffload enable the verifying and drafting to run concurrently, allowing SD to be seamlessly embedded into the offloading pipeline, while better utilizing GPU resources.
@@ 3 System Overview
During the offline phase, the target and draft models are deployed across heterogeneous memory hierarchy. SpecOffload automatically evaluates the hardware ecosystem, measuring CPU and GPU memory capacities, computational performance of CPU and GPU cores, and the bandwidth of data transfer channels between them. These hardware specifications, along with the configurations of models (➀), are input to the Adaptive Tensor Placement (§ 4.2) to determine the initial parameter allocation strategy (➁).
@@ 4.1 Interleaved Batch Pipeline
Motivated by the differing computational demands of the prefill and decoding phases in LLM inference, we introduce the Interleaved Batch Pipeline—a phase-specific pipeline design. During the prefill stage, the target model computation dominates the GPU runtime and we perform additional memory management before completion. In the decoding stage, we embed speculative decoding into the pipeline by finely interleaving the computations of the two models to fully utilize the GPU cores.
@@ 4.1.1 Prefill Phase
While our pipeline design for the prefill phase is inspired by the "zig-zag" strategy proposed by FlexGen Flexgen , we extend this approach by tailoring the micro-batch scheduling and parameter management to better support speculative decoding. In order to minimize the GPU memory footprint of the target model during the offloading stage, at the end of the prefill phase, we offload partial model parameters and the entire KV cache to CPU memory.
@@ 4.1.2 Decoding Phase
From model-level, the decoding phase involves two batches being processed alternately by the target and draft model. Conventional speculative decoding adopts a single-batch Draft-then-Verify paradigm, as the computations of the draft and target models must be executed sequentially. Consequently, the GPU resources left idle during the target model’s verification stage still remain unused, while the system suffers from additional overhead caused by switching between models.
To overcome these limitations, we propose a dual-batch interleaved design that enables true model-level parallelism. As shown on the left side of Figure 4, in time slot $t_{n}$, while the target model verifies Batch 1, the draft model concurrently generates speculative tokens for Batch 0 ($\hat{w}_{1}$, $\hat{w}_{2}$, $\hat{w}_{3}$, $\hat{w}_{4}$). Once both tasks are completed, the roles switch. In time slot $t_{n+1}$, the target model validates Batch 0 (${w}_{1}$, ${w}_{2}$, ${w}_{3}$, ${w}_{4}$), and the draft model proceeds with Batch 1. This alternating batch rotation continues until the generation is complete. This batch interleaving primarily enables parallel execution of drafting and verification.
From computation-level, the decoding pipeline involves coordination among three threads: GPU computation, CPU computation, and I/O. The right side of Figure 4 provides an illustrative example. Each layer of the target model is fine-grainedly partitioned into attention and FFN components. For clarity, lightweight components such as normalization layers are omitted from the figure, as their parameter size and computational cost are negligible. As illustrated, for the majority of the inference time, the CPU performs attention computation (current batch), the system transfers same layer’s FFN parameters from CPU memory to GPU memory (current batch), and the GPU executes draft model computation (another batch), all in parallel. Upon completion of attention computation on the CPU, the intermediate activations are transferred to the GPU. Finally, once both the parameters and activations are available on the GPU, the remaining computations are quickly completed. After completing all layers of the target model (the end of a round of parallel drafting and verification at model level), the target and draft models exchange their current batches.
@@ 5.1 Experimental Setup
Hardware. We evaluate SpecOffload in two different evironments, as shown in Table 2.
@@ 5.3 Effectiveness Analysis
We show the runtime breakdown of Mixtral 8x7B in Env #1 and 8x22B in Env #2 in Table 3 on SummEval dataset. We disable overlapping and profile the time used for GPU, I/O and CPU. Results show that our method effectively overlaps compute and I/O.
@@ A.2 Implementation
As shown in Figure 10, inter-process communication is centered around two shared memory regions, with tokens as the primary data exchanged. The Draft process writes generated draft tokens to the output shared memory for consumption by the Main process, and reads verified tokens from the input shared memory region. When a thread in the Main process reaches the verify stage and requires draft tokens, it waits for the draft model to signal availability, then read from shared memory. Upon finishing computation, the draft model signals completion and readiness for the next task using event flags, ensuring proper synchronization and preventing data races or overlap between batches. This tightly coordinated mechanism enables efficient and orderly pipelined execution across model components, while keeping both memory footprint and runtime overhead to a minimum.

### v2

@@ Abstract
Efficient LLM inference on resource-constrained devices presents significant challenges in compute and memory utilization. Due to limited GPU memory, existing systems offload model weights to CPU memory, incurring substantial I/O overhead between the CPU and GPU. This leads to two major inefficiencies: (1) GPU cores are underutilized, often remaining idle while waiting for data to be loaded; and (2) GPU memory has a low impact on performance, as reducing its capacity has minimal effect on overall throughput. In this paper, we propose SpecOffload, a high-throughput inference engine that embeds speculative decoding into offloading. Our key idea is to unlock latent GPU resources for storing and executing a draft model used for speculative decoding, thus accelerating inference at near-zero additional cost. To support this, we carefully orchestrate the interleaved execution of target and draft models in speculative decoding within the offloading pipeline, and propose a planner to manage tensor placement and select optimal parameters. Compared with the best baseline, SpecOffload improves GPU core utilization by 4.49× and boosts inference throughput by 2.54×. Our code is available at https://github.com/MobiSense/SpecOffload.
@@ 1 Introduction
As Large Language Models (LLMs) evolve, their real-world use extends far beyond chatbots to diverse applications including synthetic data generation datageneration , form processing formprocessing , and data wrangling datawrangling . These tasks are characterised by LLMs conducting offline inference in batches over a large number of tokens. For instance, corporations need to process all archives of financial documentation, whilst individuals want to construct knowledge repositories from accumulated materials. In such workloads, higher inference throughput (the number of tokens generated divided by total generation time, token/s) translates into lower total completion time, hence it is the key metric.
Privacy and cost concerns drive these tasks toward LLM deployment on edge servers or PCs, where GPU memory is a major constraint. For instance, conducting inference exclusively using GPU memory for Mixtral 8x22B (282GB) Mixtral8x22b requires at least four NVIDIA H100 (80GB) GPUs. Offloading is one of the mainstream solutions to memory-constrained inference, transferring most model parameters to more economical, capacious CPU memory and reloading them to GPU memory only when computation demands. There are also methods to overcome the memory bottleneck by compressing the model and KV cache, such as quantization, pruning, sparsification quantization ; pruning ; sparsification , etc., which are orthogonal and can be applied on top of offloading. Our focus is on designing efficient offloading strategies for high-throughput inference on resource-constrained devices.
@@
  • Underutilization of GPU cores. As shown in Figure 2, during the decoding phase, the average GPU core utilization of existing methods is only 13% at most. This inefficiency stems from GPU cores frequently remaining long time idle while awaiting parameter loading. To alleviate this issue, existing methods increase the batch size to amortize I/O overhead by loading model parameters once and reusing them across the enlarged batch, thereby improving throughput Flexgen ; Klotski . However, due to the limitation of GPU memory capacity or CPU computational capabilities, the scalability of the batch size remains inherently limited. The maximum batch size achieved by the SOTA scheme in Figure 2 is only 64, insufficient to bridge the huge gap between I/O and GPU latency (even with this batch size, the gap remains over 10×).
  • Marginal utility of GPU memory. When models far exceed GPU memory capacity, throughput remains almost unchanged even when memory usage is greatly reduced during decoding phase. We conducted experiments with FlexGen Flexgen , the most effective solution among SOTA approaches. FlexGen offloads attention computations to CPU while computing FFN layers on GPU during the decoding phase. As illustrated in Figure 2, reducing memory usage by over 5.42× for Mixtral 8×7B models led to only a 13% drop in throughput; similarly, a 2.89× reduction for Mixtral 8×22B models caused just a 5% decline. This phenomenon arises because the total volume of data to be loaded remains nearly constant, as the model size vastly exceeds the available GPU memory, leaving little room for variation even when accounting for partical parameters that can reside permanently in memory. For instance, generating a single token necessitates loading all (56) FFN layers for the leftmost blue point in Figure 2, compared to 53 layers for the rightmost blue point. Consequently, memory reduction precipitates only negligible changes in total inference time and throughput.
To harness GPU compute and memory resources more efficiently, we design SpecOffload, a novel offloading framework that unlocks latent GPU capacity by leveraging speculative decoding (SD). SD is a technique that accelerates generation by employing an auxiliary lightweight draft model to produce multiple candidate tokens, which are subsequently verified in parallel by the target model, enabling the generation of multiple tokens per forward pass blockwise ; FastInference ; ACloseLook .
@@
  • Computing draft model during GPU core idleness: SD requires the draft model to generate multiple candidate tokens in advance. Given the substantial idle periods prevalent in existing frameworks, these intervals can be utilized for completing the draft model’s computational tasks.
  • Storing draft model uses "low-yield" GPU memory: SD requires loading a draft model into memory for draft generation. We can repurpose "low-yield" memory allocations from existing frameworks to store draft model parameters and its caches instead. For instance, as shown in Figure 2, extracting 17GB of "low-yield" memory allocation provides sufficient capacity for a draft model such as Mistral 7B mistral7b to operate normally within the GPU at a small batch.
@@
  • We conduct a quantitative analysis of GPU resource utilization in representative scenarios and identify key limitations in SOTA frameworks—underutilization of GPU cores and marginal utility of GPU memory, thus reveal a novel perspective for enhancing offloading performance.
  • By designing a sub-layer model decomposition and fine-grained scheduling of compute and memory resources, we delicately embedded SD into offloading with virtually zero overhead, thereby increasing GPU core utilization by 4.49 times.
@@ 2.1 Generative LLM Inference
Large language models consist of stacked Transformer layers attentionisallyouneed . During inference, tokens are generated in an autoregressive paradigm: the prefill phase processes the complete input sequence to construct the KV cache, whereas in the decoding phase, subsequent token is produced based on the previously generated tokens and the cached KV states pageattention ; pope_efficiently_2023 ; GPT3 .
@@ 2.2 Speculative Decoding
Speculative decoding (SD) is a method for accelerating LLM inference. It adheres to a “Draft-then-Verify” framework: at each decoding step, a lightweight draft model initially proposes multiple candidate tokens (e.g., $(\hat{w}_{1},\hat{w}_{2},\hat{w}_{3},\hat{w}_{4}$)), which are collectively verified by a larger target model in a single forward pass. Only the valid subset $({w}_{1},{w}_{2})$ is accepted, after which the target model resumes decoding by independently generating the subsequent token $w_{3}$ blockwise ; FastInference ; ACloseLook . This approach enables the target model to generate multiple tokens per inference step. To further enhance the efficiency of SD, prior research has predominantly explored two avenues: the design of more effective draft models distillspec ; draft-verify ; onlineSD and draft structures medusa ; specinfer ; specexec . However, traditional single-batch SD is not well-suited for integration with offloading, as the computations of the draft and target models must be executed sequentially. As a result, the GPU resources remain underutilized during the target model’s verification phase. By introducing a dual-batch rotation strategy, SpecOffload enables the verifying and drafting to run concurrently, allowing SD to be seamlessly embedded into the offloading pipeline, while better utilizing GPU resources.
@@ 3 System Overview
During the offline phase, the target and draft models are deployed across a heterogeneous memory hierarchy. SpecOffload automatically evaluates the hardware ecosystem, measuring CPU and GPU memory capacities, computational performance of CPU and GPU cores, and the bandwidth of data transfer channels between them. These hardware specifications, along with the configurations of models (➀), are input to the Adaptive Tensor Placement (§ 4.2) to determine the initial parameter allocation strategy (➁).
@@ 4.1 Interleaved Batch Pipeline
Motivated by the differing computational demands of the prefill and decoding phases in LLM inference, we introduce the Interleaved Batch Pipeline—a phase-specific pipeline design. During the prefill stage, the target model computation dominates the GPU runtime, and we perform additional memory management before completion. In the decoding stage, we embed speculative decoding into the pipeline by finely interleaving the computations of the two models to fully utilize the GPU cores.
@@ 4.1.1 Prefill Phase
While our pipeline design for the prefill phase is inspired by the "zig-zag" strategy proposed by FlexGen Flexgen , we extend this approach by tailoring the micro-batch scheduling and parameter management to better support speculative decoding. To minimize the GPU memory footprint of the target model during the offloading stage, at the end of the prefill phase, we offload partial model parameters and the entire KV cache to CPU memory.
@@ 4.1.2 Decoding Phase
At model-level, the decoding phase involves two batches being processed alternately by the target and draft model. Conventional speculative decoding adopts a single-batch Draft-then-Verify paradigm, as the computations of the draft and target models must be executed sequentially. Consequently, the GPU resources left idle during the target model’s verification stage remain unused, while the system suffers from additional overhead caused by switching between models.
To overcome these limitations, we propose a dual-batch interleaved design that enables true model-level parallelism. As shown on the left side of Figure 4, in time slot $t_{n}$, while the target model verifies Batch 1, the draft model concurrently generates speculative tokens for Batch 0 ($\hat{w}_{1}$, $\hat{w}_{2}$, $\hat{w}_{3}$, $\hat{w}_{4}$). Once both tasks are completed, the roles switch. In time slot $t_{n+1}$, the target model validates Batch 0 (${w}_{1}$, ${w}_{2}$, ${w}_{3}$, ${w}_{4}$), and the draft model proceeds with Batch 1. This alternating batch rotation continues until the generation is complete. This batch interleaving primarily enables the parallel execution of drafting and verification.
At computation-level, the decoding pipeline involves coordination among three threads: GPU computation, CPU computation, and I/O. The right side of Figure 4 provides an illustrative example. Each layer of the target model is fine-grainedly partitioned into attention and FFN components. For clarity, lightweight components such as normalization layers are omitted from the figure, as their parameter size and computational cost are negligible. As illustrated, for the majority of the inference time, the CPU performs attention computation (current batch), the system transfers same layer’s FFN parameters from CPU memory to GPU memory (current batch), and the GPU executes draft model computation (another batch), all in parallel. Upon completion of attention computation on the CPU, the intermediate activations are transferred to the GPU. Finally, once both the parameters and activations are available on the GPU, the remaining computations are quickly completed. After completing all layers of the target model (the end of a round of parallel drafting and verification at model-level), the target and draft models exchange their current batches.
@@ 5.1 Experimental Setup
Hardware. We evaluate SpecOffload in two different evironments, as shown in Table 2. Env #2 refers to a cloud-based server.
@@ 5.3 Effectiveness Analysis
We show the runtime breakdown of Mixtral 8x7B in Env #1 and 8x22B in Env #2 in Table 3 on SummEval dataset. We disable overlapping and profile the time used for GPU, I/O, and CPU. Results show that our method effectively overlaps compute and I/O.
@@ A.2 Implementation
As shown in Figure 10, inter-process communication is centered around two shared memory regions, with tokens as the primary data exchanged. The Draft process writes generated draft tokens to the output shared memory for consumption by the Main process, and reads verified tokens from the input shared memory region. When a thread in the Main process reaches the verify stage and requires draft tokens, it waits for the draft model to signal availability, then reads from shared memory. Upon finishing computation, the draft model signals completion and readiness for the next task using event flags, ensuring proper synchronization and preventing data races or overlap between batches. This tightly coordinated mechanism enables efficient and orderly pipelined execution across model components, while keeping both memory footprint and runtime overhead to a minimum.